Projection Close 3/22
To do this evening/ was completed: Relatively quiet night on my end.
Presentation Issues:
Preshow Clips/Playlist: Double checked the playlists that are currently in use and 3/28.
Still waiting on 5 min Class preshow
Trailer Updates:
Special Events/ AV set up: Put the collection of ADA seats back in place after the third round screening of Snow White in T2
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): New DCP of REC finished transcoding and is ingesting to the LMS.
Trailer Updates:
Special Events/ AV set up: Put the collection of ADA seats back in place after the third round screening of Snow White in T2
Ordering/Receiving (posters & trailers):
Ordering/Receiving (content): New DCP of REC finished transcoding and is ingesting to the LMS.
Hanna DCP drive will (or should) arrive this Wednesday.
Outgoing/Shipping (content): Went through out current content inventory.
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning: Trash and recycling out in 4/6. New bags in.
Outgoing/Shipping (content): Went through out current content inventory.
Print/video Inspection:
Asset tracking doc updated/checked?:
Maintenance/cleaning: Trash and recycling out in 4/6. New bags in.
Cleaned the floors in 1 and 2.
Cue sheets updated/checked (every shift): Yes
Marquee updated?: Yes
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Cue sheets updated/checked (every shift): Yes
Marquee updated?: Yes
Posters updated?:
Cleaning lights on for the night?:
Projection cart by entrance?:
Additional Notes:
Comments
Post a Comment